# Inference on many jumps in nonparametric panel regression models

### v2

@@ Appendix D Additional Simulation Results
Table A1: Empirical size and power results for the uniform test for heterogeneity for DGP1–3.
@@ 3. DGP6:
We study the size and power properties of the test for the existence of jump effects for DGP4 – DGP6 in Table A2. These DGPs are the ones with the strong cross-sectional dependence we consider in our simulation study. For DGP4, we observe an empirical size slightly below the nominal size for the high-dimensional setting with $N=100$ which is due to the increase in the cross-sectional dependence in the running variable and the error term. The size properties for DGP5 are almost perfect, while the inclusion of an additional time effect in DGP6 increases the empirical size slightly. This time effect adds more noise to the system, so it is not surprising that the empirical power of DGP6 is lower than that of the other DGPs we consider.
@@
Table A2: Empirical size and power results for the uniform test of the existence of jump effects for DGP4–6.
While the preceding results are concerned with the known threshold case, we now exam the size and power properties of our uniform testing procedure in the case of unknown thresholds. We consider the same model specification as in the known threshold case. Therefore, the true threshold is $c_{0j}=0$ for all $j$. For the practical implementation of our test procedure, we consider grids of threshold locations, $[-0.6,-0.3,0,0.3,0.6]$ for DGP1 and $[-0.3,-0.15,0,0.15,0.3]$ for DGP2 and 3. The results in Table A3 reveal that the test has still good size properties, with the empirical size being smaller but close to the nominal one. As expected, the power of the test decreased compared with the known threshold case.
@@
Finally, we check the estimation accuracy of $\hat{c}_{j}$, our estimator of the unknown threshold location, $c_{0j}$. For this purpose, we set all $\gamma_{j}=T^{-2/5}(\mathrm{log}N)^{1/2}5B_{j}$, where $B_{j}\sim i.i.d.\ U[2,10]$. We consider a grid of possible threshold locations, i.e. $[-0.3,-0.29,\ldots,0.3]$. The results in Table A4 reveal that the maximum and averaged error can be reduced effectively with increasing $T$.
@@
Appendix E Empirical Applications
E.1 Application I: Jumps in the News Impact Curve
In this application, we revisit the news impact curve introduced by Engle and Ng, (1993). In particular, we want to identify possible jumps in the relationship between news events in the form of past returns on the volatility of stocks. The application serves as an illustration of our uniform testing procedure when the threshold $c_{0j}$ is unknown. Originally proposed in the context of autoregressive volatility models (e.g., ARCH), Engle and Ng, (1993) and Linton and Mammen, (2005) consider partially nonparametric specifications for the news impact curve which are able to capture possible asymmetries. A limitation of these specifications is the restriction to continuous functions. However, Zakoian, (1994) and Fornari and Mele, (1997) advocate for (parametric) GARCH specifications that allow for discontinuities in the way past returns impact volatility. We therefore want to investigate this issue by testing for the presence of jumps. In our model setup, the dependent variable $Y_{jt}$ is the Garman-Klass volatility of stock $j$ at time $t$, and $X_{jt}$ is the lagged stock return.
$Y_{jt}=f_{j}(X_{jt})+\gamma_{j}\textbf{1}_{\{X_{jt}\geq c_{0j}\}}+e_{jt}.$
We want to test the null hypothesis $H_{0}^{(1)}:\gamma_{i}=0$ for all $i\in[N]$ and all $c_{0j}$, versus the alternative $H_{a}^{(1)}:\gamma_{i}\neq 0$ for some $i$ and some $c_{0j}$. We want to reiterate that our testing procedure is robust to the presence of unobserved risk factors, $U_{jt}$, which are more than likely to occur in our application. Our data includes the daily stock returns and volatility data of the S&P 500 constituents from January 2010 to May 2024, which we obtained from Kaggle  [^1] https://www.kaggle.com/datasets/andrewmvd/sp-500-stocks. Therefore we have a fairly high-dimensional setting with $N=500$ and $T_{j}$ is the number of days of stock $j$ being a member of the stock index. As candidates for the threshold location, we consider $c_{i}\in\{-0.01,-0.005,0,0.005,0.01\}$.
The main findings are summarized in Table A5. We find significant results for $15$ out of the $500$ stocks we consider. For stocks with significant threshold effects, the locations of the estimated thresholds vary, and none of the significant threshold effects occur precisely at zero. This contradicts the rationale of the sign-switching ARCH model of Fornari and Mele, (1997), which models a jump effect based on the sign of the lagged return, and thus expects a significant threshold effect at zero. For $c_{i}=-0.01$, we get significant negative effects at the $1\%$ level for five stocks, including Allstate Corporation and Ecolab. On the other hand, we obtain significant positive effects for the $c_{i}=0.01$ threshold for eight stocks, e.g. AT&T, Booking, and Costco. These results reaffirm the findings of Chen and Ghysels, (2011), who find that both very bad news and very good news have an increasing effect on volatility. Additionally, we get another negatively significant coefficient at the $-0.005$ threshold location and another positively significant coefficient at $c_{i}=0.005$. Overall, we get evidence for the presence of jump effects in the news impact curve.
@@ E.1 Application I: Jumps in the News Impact Curve
ALL –0.01 –0.300 0.059 –5.081
ECL –0.01 –0.224 0.042 –5.374
FI –0.01 –0.217 0.044 –4.958
SJM –0.01 –0.210 0.042 –5.018
WY –0.01 –0.245 0.044 –5.511
NXPI –0.005 –0.534 0.071 –7.476
GE 0.005 0.210 0.044 4.801
T 0.01 0.171 0.032 5.357
BKNG 0.01 0.210 0.043 4.903
COST 0.01 0.154 0.030 5.102
HIG 0.01 0.249 0.048 5.224
MCK 0.01 0.232 0.050 4.641
ROST 0.01 0.209 0.040 5.250
WRB 0.01 0.161 0.032 4.971
WM 0.01 0.192 0.038 5.028
$\hat{\mathcal{I}}$ 7.476
Table A5: Uniform test for the existence of jump effects. This table reports the selected threshold locations, the estimated coefficients, the standard errors, the corresponding individual test statistic, the number of observations and the effective number of observations. Only the significant results at the $1\%$ level are reported. The last line reports the test statistic of the uniform testing procedure, with critical values $4.611$ at $1\%$ level.
E.2 Application II: Party Incumbency Effects on U.S. House Elections
In the second application, we are interested in estimating the advantage a candidate has in the U.S. House elections when the seat is currently occupied by the party of the candidate. This is called the party incumbency effect. We follow the research design of Lee, (2008). See Caughey and Sekhon, (2011) for an analysis of more recent elections. The dependent variable $Y_{jt}$ is the democratic two-party vote share in year $j$ in district $t$. The coavariate $X_{jt}$ is the difference in the two-party vote share in the previous election. Since the winner of the election is deterministically determined by the 0 threshold, our model setup is appropriate.
$Y_{jt}=f_{j}(X_{jt})+\gamma_{j}\mathbf{1}_{\{X_{jt}>0\}}+e_{jt}.$
We want to test the one-sided null hypothesis $H_{0}^{(1)}:\gamma_{i}<0$ for all $i\in[N]$, versus the alternative hypothesis $H_{a}^{(1)}$: $\gamma_{i}\geq 0$ for some $i$. Unlike Lee, (2008) who pools the data together to estimate a single incumbency effect, we conduct a separate analysis for different election years and also a separate analysis for different states by interchanging the roles of $j$ and $t$ and allowing for possible heterogeneity of incumbency effects along either the $j$ or $t$ dimension. In addition, it is easy to see our theories extend to the unbalanced data in which case we can use $T_{j}$ to denote the number of observations associated with individual/group $j$.
We use the data provided by the MIT Election Data and Science Lab, (2017). The data include the vote shares of all candidates in the elections for the US House of Representatives from 1976–2020. Due to redistricting at the start of each decade, we have to exclude years ending with a ‘0’ or ‘2’.
[…]
Now, we are interested in the state-specific effects. We restrict our attention to states with a sufficient number of elections held in the sample. We choose a threshold of 50 elections. This gives us $N=29$ states in our analysis, and $T_{j}$ now denotes the number of elections included for the $j$th state. We argue that this setting constitutes a fairly high-dimensional setting since the effective number of observations is smaller than $N$ for a fairly large proportion of the states. The results are reported in Table A8. Consistent with the previous results, almost all of the estimated coefficients are positive. By using the critical values for the simultaneous tests, we find significant effects for seven states, for Missouri and California at $1\%$ level, and for Pennsylvania, Illinois, Ohio, Iowa, and South Carolina at $5\%$. Again, we can reject the null hypothesis of no jump effect for our uniform testing procedure at a $1\%$ level. The significant states tend to belong to those with the largest number of observations. However, for the other two largest states, New York and Texas, the estimated effect is rather small and far from being significant. The incumbency effect seems to be stronger in some states compared to others. We also run the test for the homogeneity of incumbency effects ($H_{0}^{\left(2\right)}$) over the 29 states. The results are displayed in Table A9 and the test statistic is given by $\hat{\mathcal{Q}}=5.159$, which implies that we can reject the null hypothesis of homogeneous effects at the $1\%$ level. We find heterogeneity in the effects across election periods and states.
@@
1978 0.174 0.039 4.456 412 127
1984 0.163 0.028 5.826 413 154
1986 0.098 0.039 2.541 408 86
1988 0.161 0.048 3.327 395 79
1994 0.117 0.024 4.810 387 102
1996 0.037 0.022 1.645 394 112
1998 0.112 0.033 3.340 369 132
2004 0.045 0.071 0.633 359 65
2006 0.063 0.077 0.822 403 102
2008 0.063 0.023 2.788 405 142
2014 0.075 0.033 2.238 380 103
2016 0.036 0.046 0.798 386 42
2018 0.037 0.026 1.420 403 37
$\hat{\mathcal{I}}$ 5.826
Table A6: Uniform test for the existence of jump effects. This table reports the estimated coefficients, the standard errors, the corresponding individual test statistic, the number of observations and the effective number of observations. The last line reports the test statistic of the uniform testing procedure, where the critical values at $10\%$, $5\%$ and $1\%$ levels are $2.405$, $2.657$ and $3.165$, respectively.
@@
1978 0.083 0.038 2.630 412 127
1984 0.072 0.028 3.123 413 154
1986 0.008 0.038 0.631 408 86
1988 0.070 0.046 1.869 395 79
1994 0.026 0.025 1.661 387 102
1996 –0.054 0.024 –1.596 394 112
1998 0.021 0.033 1.122 369 132
2004 –0.046 0.066 –0.451 359 65
2006 –0.028 0.072 –0.160 403 102
2008 –0.028 0.024 –0.498 405 142
2014 –0.016 0.033 0.000 380 103
2016 –0.054 0.044 –0.876 386 42
2018 –0.054 0.027 –1.396 403 37
$\hat{\mathcal{Q}}$ 2.553 3.123
Table A7: Uniform test for heterogeneity. This table reports the estimated coefficients, the standard errors, the corresponding individual test statistic, the number of observations and the effective number of observations. The last line reports the test statistic of the uniform testing procedure. Critical values at $10\%$, $5\%$ and $1\%$ level are $2.405$, $2.657$ and $3.165$.
@@
CONNECTICUT 1 0.142 0.054 2.645 72 41
MASSACHUSETTS 3 0.251 0.093 2.713 129 22
NEW JERSEY 12 0.092 0.044 2.082 171 55
NEW YORK 13 0.053 0.032 1.690 387 153
PENNSYLVANIA 14 0.096 0.032 2.984 237 100
ILLINOIS 21 0.114 0.039 2.922 250 85
INDIANA 22 0.080 0.029 2.720 125 62
MICHIGAN 23 0.067 0.036 1.874 206 79
OHIO 24 0.133 0.045 2.977 230 69
WISCONSIN 25 0.128 0.075 1.704 97 31
IOWA 31 0.125 0.040 3.154 64 50
MISSOURI 34 0.258 0.029 8.878 115 37
VIRGINIA 40 0.223 0.102 2.175 102 54
ALABAMA 41 0.212 0.070 3.020 74 21
FLORIDA 43 0.039 0.049 0.798 220 85
GEORGIA 44 0.101 0.048 2.101 126 40
LOUISIANA 45 –0.403 0.228 –1.764 80 16
MISSISSIPPI 46 0.101 0.095 1.057 51 13
NORTH CAROLINA 47 0.027 0.040 0.673 147 77
SOUTH CAROLINA 48 0.352 0.108 3.249 65 25
TEXAS 49 0.016 0.067 0.246 344 117
KENTUCKY 51 0.067 0.057 1.186 70 33
MARYLAND 52 0.126 0.134 0.937 105 29
TENNESSEE 54 0.049 0.059 0.828 101 20
ARIZONA 61 0.131 0.057 2.275 75 34
COLORADO 62 0.070 0.059 1.200 78 44
CALIFORNIA 71 0.088 0.025 3.458 639 189
OREGON 72 0.046 0.047 0.970 64 33
WASHINGTON 73 0.036 0.041 0.878 114 55
$\hat{\mathcal{I}}$ 8.878
Table A8: Uniform test for the existence of jump effects. This table reports the estimated coefficients, the standard errors, the corresponding individual test statistic, the number of observations and the effective number of observations. The last line reports the test statistic of the uniform testing procedure, where the critical values at $10\%$, $5\%$ and $1\%$ levels are $2.685$, $2.917$ and $3.392,$ respectively. Due to the low effective sample size, the bandwidth is chosen uniformly over all states.
@@
CONNECTICUT 1 0.046 0.054 0.854 72 41
MASSACHUSETTS 3 0.155 0.090 1.714 129 22
NEW JERSEY 12 –0.004 0.045 –0.085 171 55
NEW YORK 13 –0.043 0.034 –1.271 387 153
PENNSYLVANIA 14 0.000 0.034 0.000 237 100
ILLINOIS 21 0.017 0.040 0.436 250 85
INDIANA 22 –0.016 0.032 –0.512 125 62
MICHIGAN 23 –0.029 0.037 –0.769 206 79
OHIO 24 0.037 0.045 0.810 230 69
WISCONSIN 25 0.031 0.074 0.427 97 31
IOWA 31 0.028 0.041 0.701 64 50
MISSOURI 34 0.161 0.031 5.159 115 37
VIRGINIA 40 0.127 0.100 1.268 102 54
ALABAMA 41 0.116 0.069 1.673 74 21
FLORIDA 43 –0.057 0.049 –1.157 220 85
GEORGIA 44 0.005 0.048 0.099 126 40
LOUISIANA 45 –0.499 0.221 –2.261 80 16
MISSISSIPPI 46 0.005 0.093 0.051 51 13
NORTH CAROLINA 47 –0.069 0.041 –1.682 147 77
SOUTH CAROLINA 48 0.256 0.106 2.427 65 25
TEXAS 49 –0.080 0.066 –1.215 344 117
KENTUCKY 51 –0.029 0.056 –0.513 70 33
MARYLAND 52 0.029 0.130 0.226 105 29
TENNESSEE 54 –0.048 0.058 –0.814 101 20
ARIZONA 61 0.034 0.057 0.603 75 34
COLORADO 62 –0.026 0.058 –0.446 78 44
CALIFORNIA 71 –0.009 0.028 –0.307 639 189
OREGON 72 –0.050 0.048 –1.048 64 33
WASHINGTON 73 –0.060 0.042 –1.431 114 55
$\hat{\mathcal{Q}}$ 5.159
@@
Appendix F Key Intuition on the Dependence Structure

### v3

@@ E.2 Additional Simulation Results
Table A2: Empirical size and power results for the uniform test of heterogeneity for DGP1–3.
@@ 3. DGP6:
We study the size and power properties of the test for the existence of jump effects for DGP4 – DGP6 in Table A3. These DGPs are the ones with the strong cross-sectional dependence we consider in our simulation study. For DGP4, we observe an empirical size slightly below the nominal size for the high-dimensional setting with $N=100$ which is due to the increase in the cross-sectional dependence in the covariate and the error term. The sizes for DGP5 are almost perfect, while the inclusion of an additional time effect in DGP6 increases the empirical size slightly. This time effect adds more noise to the system, so it is not surprising that the empirical power of DGP6 is lower than that of the other DGPs we consider.
@@
Table A3: Empirical results on the size and power of the uniform test for the existence of jump effects for DGP4–6.
While the preceding results are concerned with the known threshold case, we now exam the size and power properties of our uniform testing procedure in the case of unknown thresholds. We consider the same model specification as in the known threshold case. Therefore, the true threshold is $c_{0j}=0$ for all $j$. For the practical implementation of our test procedure, we consider the following grids of threshold locations. Due to different ranges in the covariate distribution, we consider the grid $[-0.6,-0.3,0,0.3,0.6]$ for DGP1 and $[-0.3,-0.15,0,0.15,0.3]$ for DGP2 and 3. The results in Table A4 reveal that the test has still good size properties, with the empirical size being slightly smaller than nominal in some settings and slightly larger in others. As expected, the power of the test decreases compared to the case with a known threshold. However, as $T$ increases, the power still approaches one.
@@
Apart from the size and power properties of our uniform testing procedure, we also check the estimation accuracy of $\hat{c}_{j}$ ( estimator of the unknown threshold location $c_{0j}$). For this purpose, we set all $\gamma_{j}=T^{-2/5}(\mathrm{log}N)^{1/2}5B_{j}$, where $B_{j}\sim i.i.d.\ U[2,10]$. We consider a grid of possible threshold locations, i.e. $[-0.3,-0.29,\ldots,0.3]$. The results in Table A5 reveal that the maximum and averaged error over $j$ can be reduced effectively with increasing $T$.
@@
In the following, we compare our uniform testing procedure to that of a pooled parametric threshold model. To be precise, as a comparison we assume the following linear panel threshold model (Hansen, , 2000) (see also Section G.3 for details),
$Y_{jt}=\beta_{0}+\beta_{1}X_{jt}+\mathbf{1}_{\{X_{jt}\geq c_{0}\}}\left(\beta_{0}^{+}+\beta_{1}^{+}X_{jt}\right)+e_{jt}.$
The parameter of interest is $\gamma:=\beta_{0}^{+}+\beta_{1}^{+}c_{0}$ and we test the null hypothesis $H_{0}:\gamma=0$. We consider sup-Wald type statistics discussed by Andrews, (1993). For this comparison, we consider the following linear data generating process.
  1. DGP7:
    Linear model. The smooth part of the conditional mean function is now linear.
    $Y_{jt}=0.2X_{jt}+\gamma_{j}\mathbf{1}_{\{X_{jt}\geq c_{0j}\}}+\sigma(X_{jt},U_{jt})\epsilon_{jt}.$
    The generation of $X_{jt}$, $U_{jt}$, and $\sigma(\cdot,\cdot)$ is identical to DGP3. A proportion between 00 and $1$ of the coefficients $\gamma_{j}$ are drawn from a standard normal distribution, the remaining coefficients are set to zero.
The comparison results are presented in Table A6. The results confirm that our uniform testing procedure has similar size properties as a parametric test but has much larger power under the alternative. This is true for the sparse case with a proportion of only $10\%$ non-zero signals, as well as for the relatively dense case with a proportion of $20\%$. The reasons for this are signal-dilution on the one hand and signal cancellation on the other hand. Regarding the former, our uniform testing procedure is robust towards settings with sparse signal, in which a pooled test suffers from dilution of signals. And regarding the latter, our uniform testing procedure does not have the problem of signal cancellation in case the signals for different $j$’s have opposite signs and cancel each other out.
Uniform test Pooled parametric
Proportion $N$ $T$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$
0% 50 400 0.085 0.036 0.007 0.126 0.080 0.017
50 800 0.117 0.054 0.013 0.106 0.058 0.011
100 400 0.123 0.055 0.014 0.116 0.067 0.020
100 800 0.106 0.060 0.011 0.113 0.055 0.019
10% 50 400 0.451 0.368 0.249 0.251 0.158 0.066
50 800 0.743 0.682 0.576 0.300 0.215 0.115
100 400 0.608 0.526 0.380 0.216 0.128 0.051
100 800 0.894 0.858 0.753 0.307 0.231 0.120
20% 50 400 0.660 0.583 0.430 0.325 0.225 0.122
50 800 0.896 0.865 0.780 0.392 0.316 0.210
100 400 0.842 0.792 0.661 0.298 0.225 0.110
100 800 0.983 0.967 0.938 0.407 0.327 0.211
Table A6: Empirical size and power results for the uniform test of the existence of jump effects under unknown threshold locations for DGP7. Comparison with a sup-Wald test based on a linear panel threshold model.
We obtain similar results in the comparison with a pooled nonparametric test in Table A7. The results are based on DGP3 and the non-zero coefficients are generated from a standard normal distribution. For the comparison, we pool the observations along the individual dimension. Again, the uniform testing procedure is more powerful than the pooled testing procedure in all settings with existence of some non-zero coefficients.
Uniform test Pooled nonparametric
Proportion $N$ $T$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$ $\alpha=0.1$ $\alpha=0.05$ $\alpha=0.01$
0% 50 400 0.120 0.061 0.012 0.099 0.054 0.009
50 800 0.117 0.061 0.011 0.083 0.044 0.004
100 400 0.122 0.067 0.019 0.094 0.050 0.011
100 800 0.117 0.053 0.018 0.106 0.055 0.014
10% 50 400 0.460 0.376 0.235 0.124 0.072 0.013
50 800 0.693 0.633 0.522 0.144 0.076 0.029
100 400 0.651 0.569 0.414 0.140 0.082 0.021
100 800 0.862 0.823 0.729 0.157 0.095 0.028
20% 50 400 0.653 0.575 0.441 0.148 0.086 0.026
50 800 0.915 0.885 0.790 0.206 0.144 0.061
100 400 0.807 0.737 0.591 0.144 0.097 0.027
100 800 0.983 0.971 0.932 0.215 0.124 0.055
Table A7: Empirical size and power results for the uniform test of the existence of jump effects under unknown threshold locations for DGP3. Comparison with pooled nonparametric test.
Finally, we examine the size and power properties of the test of existence of a threshold effect in the first derivative of the conditional expectations. We restrict our simulations on the known threshold case. Under the null, we set $f_{j}(x,u)=5x$, whereas we consider $f_{j}(x,u)=5|x|$ under the alternative hypothesis. The share of individuals’ DGPs generated under the alternative is $20\%$ for both $N=10$ and $N=100$. For this setting, we oversmooth the bandwidth. The results for DGP1–DGP3 are presented in Table A8. The size properties are reasonably close the nominal size. Similarly, the power approaches one with increasing sample size and dimension.
Size Power
DGP1 10 400 0.102 0.052 0.007 0.923 0.831 0.527
10 800 0.113 0.046 0.011 0.998 0.995 0.941
100 400 0.120 0.058 0.008 0.999 0.998 0.909
100 800 0.115 0.060 0.010 1.000 1.000 1.000
DGP2 10 400 0.109 0.050 0.014 0.718 0.590 0.327
10 800 0.096 0.048 0.009 0.956 0.915 0.716
100 400 0.100 0.040 0.009 0.978 0.935 0.710
100 800 0.138 0.054 0.012 1.000 1.000 0.997
DGP3 10 400 0.108 0.064 0.014 0.710 0.564 0.303
10 800 0.128 0.067 0.013 0.946 0.903 0.703
100 400 0.130 0.054 0.010 0.988 0.939 0.693
100 800 0.126 0.057 0.016 1.000 1.000 0.999
Table A8: Empirical size and power results for the uniform test of the existence of derivative jump effects under known threshold locations for DGP1–3.
Appendix F Empirical Applications
In this section, we present further estimation results for our two empirical applications. First, we study possible jumps in the news impact curve for constituents of the S&P 500. As a second application, we study the impact of the incumbency effect in U.S. House elections.
F.1 Application I: Jumps in the News Impact Curve
@@
ALL –0.010 –0.240 0.043 –5.535
AXP –0.010 –0.223 0.048 –4.702
CTAS –0.010 –0.189 0.041 –4.632
ORLY –0.010 –0.211 0.044 –4.750
SJM –0.010 –0.191 0.040 –4.742
NXPI –0.005 –0.529 0.078 –6.793
HON 0.005 0.136 0.028 4.763
T 0.010 0.162 0.033 4.857
MCK 0.010 0.211 0.042 5.000
TRV 0.010 0.156 0.031 4.978
WM 0.010 0.181 0.034 5.290
$\hat{\mathcal{I}}$ 6.793
Table A9: Uniform test for the existence of jump effects. This table reports the selected threshold locations, the estimated coefficients, the standard errors, and the corresponding individual test statistic. Only the significant results at the $1\%$ level are reported. The last line reports the test statistic of the uniform testing procedure, with critical values $4.610$ at $1\%$ level.
F.2 Application II: Party Incumbency Effects on U.S. House Elections
@@
1978 0.174 0.043 4.103 412 127
1984 0.163 0.030 5.515 413 154
1986 0.098 0.041 2.418 408 86
1988 0.161 0.042 3.830 395 79
1994 0.117 0.035 3.325 387 102
1996 0.037 0.034 1.087 394 112
1998 0.112 0.035 3.167 369 132
2004 0.045 0.066 0.676 359 65
2006 0.063 0.052 1.222 403 102
2008 0.063 0.029 2.179 405 142
2014 0.075 0.026 2.912 380 103
2016 0.036 0.040 0.902 387 42
2018 0.037 0.032 1.171 403 37
$\hat{\mathcal{I}}$ 5.515
Table A10: Uniform test for the existence of jump effects. This table reports the estimated coefficients, the standard errors, the corresponding individual test statistic, the number of observations (Obs) and the effective number of observations (EObs). The last line reports the test statistic of the uniform testing procedure, where the critical values at $10\%$, $5\%$ and $1\%$ levels are $2.406$, $2.657$ and $3.166$, respectively.
@@
1978 0.100 0.041 2.451 412 127
1984 0.089 0.029 3.011 413 154
1986 0.024 0.039 0.606 408 86
1988 0.086 0.040 2.140 395 79
1994 0.042 0.034 1.229 387 102
1996 –0.038 0.033 –1.146 394 112
1998 0.037 0.034 1.078 369 132
2004 –0.030 0.062 –0.481 359 65
2006 –0.011 0.049 –0.234 403 102
2008 –0.012 0.029 –0.414 405 142
2014 0.000 0.026 0.000 380 103
2016 –0.038 0.039 –0.986 387 42
2018 –0.038 0.031 –1.198 403 37
$\hat{\mathcal{Q}}$ 3.011
Table A11: Uniform test for heterogeneity. This table reports the estimated coefficients, the standard errors, the corresponding individual test statistic, the number of observations and the effective number of observations. The last line reports the test statistic of the uniform testing procedure. Critical values at $10\%$, $5\%$ and $1\%$ level are $2.406$, $2.657$ and $3.166$.
@@
CONNECTICUT 1 0.142 0.048 2.949 72 41
MASSACHUSETTS 3 0.251 0.184 1.362 129 22
NEW JERSEY 12 0.092 0.047 1.953 171 55
NEW YORK 13 0.053 0.026 2.032 387 153
PENNSYLVANIA 14 0.096 0.036 2.699 237 100
ILLINOIS 21 0.114 0.043 2.658 250 85
INDIANA 22 0.080 0.031 2.546 125 62
MICHIGAN 23 0.067 0.032 2.086 206 79
OHIO 24 0.133 0.047 2.815 230 69
WISCONSIN 25 0.128 0.087 1.472 97 31
IOWA 31 0.125 0.038 3.247 64 50
MISSOURI 34 0.258 0.050 5.117 115 37
VIRGINIA 40 0.223 0.069 3.209 102 54
ALABAMA 41 0.212 0.068 3.124 74 21
FLORIDA 43 0.039 0.066 0.593 221 85
GEORGIA 44 0.101 0.075 1.340 126 40
LOUISIANA 45 –0.403 0.386 –1.044 80 16
MISSISSIPPI 46 0.101 0.121 0.836 51 13
NORTH CAROLINA 47 0.027 0.039 0.684 147 77
SOUTH CAROLINA 48 0.352 0.106 3.324 65 25
TEXAS 49 0.016 0.060 0.271 344 117
KENTUCKY 51 0.067 0.086 0.786 70 33
MARYLAND 52 0.126 0.096 1.314 105 29
TENNESSEE 54 0.049 0.128 0.381 101 20
ARIZONA 61 0.131 0.071 1.836 75 34
COLORADO 62 0.070 0.053 1.322 78 44
CALIFORNIA 71 0.088 0.027 3.252 639 189
OREGON 72 0.046 0.042 1.086 64 33
WASHINGTON 73 0.036 0.037 0.981 114 55
$\hat{\mathcal{I}}$ 5.117
Table A12: Uniform test for the existence of jump effects. This table reports the estimated coefficients, the standard errors, the corresponding individual test statistic, the number of observations and the effective number of observations. The last line reports the test statistic of the uniform testing procedure, where the critical values at $10\%$, $5\%$ and $1\%$ levels are $2.685$, $2.917$ and $3.392,$ respectively. Due to the low effective sample size, the bandwidth is chosen uniformly the same over all states.
Figure F.4: Local linear fit of a significant threshold effect for California (left panel) and an insignificant effect for Pennsylvania (right panel).
@@
CONNECTICUT 1 0.046 0.050 0.913 72 41
MASSACHUSETTS 3 0.155 0.179 0.866 129 22
NEW JERSEY 12 –0.004 0.049 –0.077 171 55
NEW YORK 13 –0.043 0.032 –1.347 387 153
PENNSYLVANIA 14 0.000 0.039 0.000 237 100
ILLINOIS 21 0.017 0.045 0.384 250 85
INDIANA 22 –0.016 0.036 –0.453 125 62
MICHIGAN 23 –0.029 0.036 –0.789 206 79
OHIO 24 0.037 0.049 0.744 230 69
WISCONSIN 25 0.031 0.086 0.366 97 31
IOWA 31 0.028 0.042 0.684 64 50
MISSOURI 34 0.161 0.052 3.096 115 37
VIRGINIA 40 0.127 0.070 1.818 102 54
ALABAMA 41 0.116 0.068 1.698 74 21
FLORIDA 43 –0.057 0.066 –0.86 221 85
GEORGIA 44 0.005 0.075 0.064 126 40
LOUISIANA 45 –0.499 0.373 –1.339 80 16
MISSISSIPPI 46 0.005 0.118 0.040 51 13
NORTH CAROLINA 47 –0.069 0.043 –1.627 147 77
SOUTH CAROLINA 48 0.256 0.104 2.463 65 25
TEXAS 49 –0.080 0.061 –1.301 344 117
KENTUCKY 51 –0.029 0.085 –0.342 70 33
MARYLAND 52 0.029 0.094 0.313 105 29
TENNESSEE 54 –0.048 0.125 –0.381 101 20
ARIZONA 61 0.034 0.071 0.484 75 34
COLORADO 62 –0.026 0.055 –0.475 78 44
CALIFORNIA 71 –0.009 0.032 –0.269 639 189
OREGON 72 –0.050 0.045 –1.114 64 33
WASHINGTON 73 –0.060 0.040 –1.496 114 55
$\hat{\mathcal{Q}}$ 3.096
@@
Appendix G Addtional Remarks
G.1 Key Intuition on the Dependence Structure